OpenStudy (anonymous):

HELP!!! Find P(not green) the probability of selecting anything but a green M&M at random from a large bag

OpenStudy (zzr0ck3r):

amen

OpenStudy (zzr0ck3r):

1-.16

OpenStudy (zzr0ck3r):

there is a 16% chance of getting green so there is a 100%-16% chance of not getting a green

OpenStudy (anonymous):

you sure lol i dont need to first find what yellow is??

OpenStudy (anonymous):

@myko help

OpenStudy (zzr0ck3r):

yellow is 1 - all the others added up

OpenStudy (zzr0ck3r):

but it does not matter to what you asked

OpenStudy (zzr0ck3r):

if p = .16 what is p^not?

OpenStudy (anonymous):

he is right, you know that the probability of all the posibilityes together is 1, so just rest the probability of the green one

OpenStudy (anonymous):

so its 0.16?

OpenStudy (anonymous):

but if you whant you could first find the yellow one, and later add all but green

OpenStudy (zzr0ck3r):

1-.16

OpenStudy (zzr0ck3r):

.16 is chance of getting green

OpenStudy (anonymous):

0.84??

OpenStudy (zzr0ck3r):

if i told you, you had a red ball a blue ball and a green ball in a bag, the chances of you getting a green are .33333 what are the chances you get a red or a blue?

OpenStudy (anonymous):

the same?

OpenStudy (anonymous):

but adding the red & blue?

OpenStudy (zzr0ck3r):

1-.333 0r yes add the red a blue, they are 1/3 each so 1/3+1/3 = 2/3 or 1-1/3 = 2/3

OpenStudy (zzr0ck3r):

probabilities must add to 1, so if the prob of getting green is .16 there is only one other choice...not green so 1-prob of getting green

OpenStudy (anonymous):

ohhhh ok i kinda get thiss

OpenStudy (anonymous):

i have another question...how do i know which probability's cant happen?

OpenStudy (anonymous):

how do i know?

OpenStudy (zzr0ck3r):

if they have a probability, then they can happen.

OpenStudy (zzr0ck3r):

so since we have all the colors adding to 1, then the prob of getting a black m&m is 0

OpenStudy (anonymous):

right but like say they give me random numbers like 1/16, 0.001, 5/4, 0.82

OpenStudy (anonymous):

how do i know which one cant happen?

OpenStudy (zzr0ck3r):

well this makes no since because they all add to over 1

OpenStudy (zzr0ck3r):

sense*

OpenStudy (zzr0ck3r):

but if there is a number then it can happen, a prob of .00000000000001 will most likely not happen, but there is a very very very small chance it will. imagine 1 M&M in a pool full of m&m's

OpenStudy (zzr0ck3r):

5/4 is > 1

OpenStudy (zzr0ck3r):

so we cant have a 120% chance of something happening.

OpenStudy (zzr0ck3r):

but this was not a set of probabilities, it was just 4 different probabilities, but ones does not make sense to talk about.

OpenStudy (anonymous):

so it would be 0.0001?

OpenStudy (zzr0ck3r):

no, it is the one that is greater than 1

OpenStudy (anonymous):

5/4?